OpenStudy (anonymous):

The base of a triangle is 8cm greater than the height. the area is 64cm^2 Height? base?

OpenStudy (anonymous):

height 2056 base 2040

OpenStudy (anonymous):

right???

OpenStudy (anonymous):

other way around base 2056 height 2040

OpenStudy (anonymous):

let the base be equal to 8xcm and the height = x, the u know the area ofa triangle is equal A=1/2(b)(h). This means 64=1/2(8x)(x), then solve for x and that will be your height, in order to get your base take that value of x and multiply by 8.

OpenStudy (anonymous):

u should get x=4, the BAse=32cm and height =4cm